OpenStudy (anonymous):

Find the seventh term of the following sequence? 1.5, 3, 6, ...

OpenStudy (anonymous):

18?

OpenStudy (anonymous):

96

OpenStudy (btaylor):

it is multiplied by 2 each time. A general equation for the nth term:\[a_n = 1.5 (2)^{n-1}\]So, for n=7, \[a_7 = 1.5 (2^6) = 1.5 \times 64 = 96\]